I don't understand the answer in part b. For Moonited Republic, 10w=4c To find the cost of 1w, divide both sides by 10 and we get 1w=.4c, but the answer says 1c=.4w. I don't understand why they flipped the numbers.Step 2 of 6 A b. Moonited Republic has a comparative advantage in cheese. The opportunity cost of producing a kilogram of che8. Consider another Ricardian example, using standard Ricardian assumptions: Labor Hours per Bottle of Wine Labor Hours per K
Answer #1

b) Comparative Advantage

A country has a comparative advantage in producing that good if the opportunity cost of producing that good is lower in that country as compared to another country.

Vintland:

15 bottles of wine = 10 kilograms of cheese

1 bottle of wine = 10/15 = 0.67 kilograms of cheese

1 kilogram of cheese = 15/10 = 1.5 bottles of wine

Opportunity cost of producing 1 bottle of wine is 0.67 kilograms of cheese and 1 kilogram of cheese is 1.5 bottles of wine.

Moonited Republic:

10 bottles of wine = 4 kilograms of cheese

1 bottle of wine = 4/10 = 0.4 kilograms of cheese

1 kilogram of cheese = 10/4 = 2.5 bottles of wine

Opportunity cost of producing 1 bottle of wine is 0.4 kilograms of cheese and 1 kilogram of cheese is 2.5 bottles of wine.

Vintland has comparative advantage in the production of cheese while Moonited Republic has comparative advantage in the production of wine.